# How to Train a Backdoor-Robust Model on a Poisoned Dataset without Auxiliary Data?

### v1

@@ V-B1 Exploring the Relationship between Perturbation and Backdoor
where $\delta=\max(\min(\frac{\nabla_{\delta}\ell(f_{\theta}(x),y)}{\|\nabla_{\delta}\ell(f_{\theta}(x),y)\|_{2}},\epsilon),-\epsilon)$ and $\hat{m}$ is a randomly-generated mask of the perturbation. As shown in Fig.5, on both BadNets and Blend attacks, most poisoned samples exhibit good consistency against adversarially generated perturbations $\delta$. To further explain this intriguing phenomenon, we propose the corresponding proof based on the recent studies [47] to analyze the characteristics of poisoned samples (detailed proof can be found in Appendix VIII):
Figure 4: Loss landscape of clean sample and poisoned sample.
Theorem 1. For $\mathcal{D}$, $\mathcal{D}_{p}$ that satisfy $N_{b}=|\mathcal{D}|$ and $N_{p}=|\mathcal{D}_{p}|$, i.i.d. sampled from an uniform distribution and belonging to $K$ classes. Assume that deep neural network $f(\cdot;\theta)$ is a multivariate kernel regression (RBF kernel) with the same objective as attackers. For a given attacked sample $\boldsymbol{x}^{\prime}=(\mathbf{1}-\boldsymbol{m})\times\boldsymbol{x}+\boldsymbol{m}\times\boldsymbol{t}$, we have: $\lim_{N_{p}\rightarrow N_{b}}C\left(\boldsymbol{x}^{\prime}+\delta\right)=y_{t}$.
@@
Figure 5: Statistical histogram of training samples.
@@
The theorem above demonstrates that when the size of $\mathcal{D}_{p}$ is close to that of the clean ones, poisoned samples present robustness towards potential perturbations. To figure it out intuitively, we also visualize the loss landscape of the clean and poisoned samples (Blend Attack), respectively, as shown in Fig.4. Note that the loss of the clean sample exhibits steeper changes when the perturbation is performed in the direction of the gradient. This observation also further confirms the previous theoretical analysis.
@@ V-B2 Dataset Partition
Algorithm 1 shows the complete process of our dataset partition operation. This algorithm delineates a meticulous process for partitioning a dataset into trusted and potentially compromised subsets by exploiting the variances in the behavior of a model when subjected to perturbed inputs. At its core, the algorithm iterates over each sample $(x,y)$ in the dataset $\mathcal{D}_{u}$, applying a strategically-sized perturbation patch characterized by a radius $r$ and bounded by a perturbation radius $\epsilon$, to the input sample. This is achieved through generating a random patch mask $\hat{m}$ and calculating an optimal adversarial perturbation $\delta$, which is applied to the input $x$ to yield a perturbed counterpart $\hat{x}$. The key metric for partitioning, the KL divergence, is computed between the model’s predictions on the original and perturbed inputs, with these divergence scores being aggregated into a list $Q$. Subsequently, the dataset is partitioned into two subsets, $\mathcal{D}_{u}$ and $\mathcal{D}_{c}$, based on the divergence scores relative to a partition rate $p$, indexing subsets with assumed clean and those potentially compromised, respectively. This method essentially leverages the sensitivity of backdoored models to perturbations in identifying and isolating suspicious data samples, thus serving as a robust strategy in the preprocessing phase of data handling. With the above steps, we obtain $\mathcal{D}_{u}$ and $\mathcal{D}_{c}$. Here, $\mathcal{D}_{c}$ denotes the selected small number of clean samples. $\mathcal{D}_{u}$ represents the rest of the training data that contains the poisoned samples.
V-B3 Backdoor Enhancement and Standard Training
@@
V-B4 Relabel and Relearn
To obtain a completed clean dataset and further improve the performance of the clean model, AdvrBD relabels the poisoned samples and merges them with clean samples to obtain a clean and complete dataset:
@@ V-B4 Relabel and Relearn
VI Experiments
In this section, we comprehensively evaluate the performance of the proposed AdvrBD. Firstly, we show the experimental settings, including the experimental environment, datasets, networks, backdoor attacks, and backdoor defenses. Then, we analyze and summarize the experimental results. Finally, we also present the ablation studies.
VI-A Experimental Settings
Experimental Environment: All experiments were conducted on a server equipped with Intel i9-9900K, 3.60GHz processor, 32GB RAM, NVIDIA GeForce RTX 3090, and PyTorch.
Datatsets: We perform our experiments on four image datasets, including Imagnette, Tiny ImageNet, CIFAR10, and CIFAR100. The introduction of the above four datasets is shown as follows.
@@ VI-A Experimental Settings
  • CIFAR10[50]: CIFAR10 dataset is a database of tiny images, including $50,000$ training images and $10,000$ testing images with $10$ classes.
@@
Networks: We perform our experiments on three networks, including ResNet18, ResNet34, and MobileNetV2. It should be noted that unless otherwise specified, the default experimental network is ResNet18.
Backdoor Attacks: We employ $8$ state-of-the-art backdoor attacks to evaluate the proposed AdvrBD in our experiments, including three dirty-label attacks: BadNet [9], Blend [19], and PhysicalBA [40], one clean-label attack: SIG [42], one input-specific-trigger attack: WaNet [41], and three adaptive attacks: TaCT [37], AdaptiveBlend [43] and AdaptivePatch [43]. The detailed configurations of these backdoor attacks are shown in Table I.
@@
Backdoor Defense: We compare AdvrBD with $7$ established backdoor defense baselines, including FineTuning [51], FinePruning [51], CutMix [52], CLP [18], DBR [13], SCAnFT [37], and ABL [12].
@@
Evaluation Metrics: We employ the attack success rate (ASR) and accuracy (ACC) to evaluate the performance of the proposed backdoor defense framework. ASR indicates the ratio of the triggered samples that are misclassified as the target label. ACC represents the accuracy of the model on the clean samples. The lower ASR and the higher ACC indicate the better performance of the backdoor defense mechanism.
@@
TABLE II: The comparison results (%) between our AdvrBD and $7$ state-of-the-art backdoor defenses methods against $5$ state-of-the-art backdoor attacks on ResNet18 with $5\%$ poisoned ratio.
No Defense FineTuning FinePruning CutMix CLP DBR SCAnFT ABL AdvrBD
@@
Average 94.14 84.28 68.09 60.72 59.28 60.94 64.76 69.31 23.21 25.76 60.94 28.42 55.13 11.41 47.46 60.49 70.20 0.00
CIFAR10 BadNet 92.00 95.20 83.30 96.81 75.05 93.19 81.21 83.30 34.18 0.55 67.47 1.21 69.89 2.09 71.98 0.00 85.72 0.00
@@
Average 87.12 88.36 83.94 92.61 79.01 78.70 80.57 78.48 32.57 30.81 40.06 41.01 70.22 13.65 82.20 28.72 85.14 0.53
CIFAR100 BadNet 72.10 98.20 58.40 91.40 42.00 91.09 42.91 86.03 54.55 83.30 29.35 0.10 56.78 0.00 63.16 0.10 60.53 0.00
@@
Average 60.62 88.36 61.92 87.21 40.81 57.18 49.15 85.99 45.32 77.99 20.41 52.68 50.12 1.01 62.50 19.15 58.99 0.22
TinyImageNet BadNet 37.80 97.95 46.00 98.54 26.87 97.58 31.55 90.84 5.08 84.90 0.70 0.00 32.21 0.00 46.80 0.00 35.73 0.00
@@
Average 38.79 98.80 48.22 98.55 25.84 83.14 33.95 96.46 9.07 51.90 4.88 46.10 24.61 0.00 38.04 55.67 37.86 0.00
VI-B AdvrBD Performance
VI-B1 Comparing to the Existing Defenses
In this paper, we employ $8$ state-of-the-art backdoor attacks (including $5$ familiar backdoor attacks and $3$ adaptive backdoor attacks) to demonstrate the effectiveness of our AdvrBD. In this section, we take into account $5$ state-of-the-art backdoor attacks and compare the performance of AdvrBD with $7$ other backdoor defense techniques. Table II shows the performance of the proposed AdvrBD method on Imagenette, CIFAR10, CIFAR100, and TinyImageNet. Obviously, our AdvrBD achieves the best results in reducing ASR against most backdoor attacks while maintaining a satisfactory ACC across all $4$ datasets.
Concretely, we mainly have five observations from Table II. The detailed analysis is shown as follows:
@@ VI-B1 Comparing to the Existing Defenses
(4) All the existing defense methods have an unstable performance against different backdoor attacks. They perform satisfactorily on some backdoor attacks and may be ineffective against the other backdoor attacks. This greatly reduces the availability of these defense methods in real-world scenarios because it is impractical for the defender to know the backdoor methods in advance. For example, ABL performs well on the BadNet attack but has an unsatisfactory performance on the SIG attack. Note that we have observed the different performance of these defense methods in different papers because their hyperparameters are different. We argue that it’s not practical to set different hyperparameters for different attacks because defenders have no knowledge about the potential attacks. Therefore, we use the same hyperparameter configurations for other defenses as well as AdvrBD.
(5) Compared with the $7$ state-of-the-art backdoor defense techniques, our AdvrBD can achieve the lowest ASR while maintaining a satisfactory performance on the clean samples. Moreover, the proposed AdvrBD has a more stable performance on different datasets and various backdoor attacks.
TABLE III: The comparison results (%) between our AdvrBD and $7$ state-of-the-art backdoor defenses methods against $3$ state-of-the-art adaptive attacks on ResNet18 with $5\%$ poisoned ratio.
No Defense FineTuning FinePruning CutMix CLP DBR SCAnFT ABL AdvrBD
@@
Average 94.60 62.83 70.91 4.99 59.00 27.70 68.43 22.53 44.04 10.34 56.88 30.65 54.94 4.43 40.16 32.92 69.99 0.28
CIFAR10 TaCT 82.10 63.20 82.63 65.71 75.05 65.71 78.68 41.43 37.91 3.08 77.58 63.19 72.42 62.20 85.16 48.02 82.42 0.00
@@
Average 83.63 51.27 82.93 60.66 79.45 40.07 71.76 32.34 52.56 3.55 67.69 33.52 74.65 20.81 85.71 50.26 82.86 0.00
CIFAR100 TaCT 55.40 22.80 58.70 25.00 42.00 17.71 53.24 17.41 18.72 32.29 41.40 29.15 44.13 0.61 63.26 0.91 61.84 0.00
@@
Average 47.93 55.47 58.23 56.24 36.47 27.56 46.22 20.82 28.81 13.19 17.14 42.85 52.87 17.81 61.14 12.58 61.61 0.00
TinyImageNet TaCT 36.95 13.65 46.55 10.87 26.07 8.05 32.36 19.38 2.62 0.00 0.55 0.00 25.42 0.00 40.71 20.73 38.80 0.00
@@
Average 36.75 37.13 46.55 61.77 24.43 25.46 32.61 43.10 9.36 9.68 0.40 25.88 27.51 0.00 37.90 57.49 39.78 0.00
VI-B2 The Resistance to Adaptive Attacks

### v2

@@ 4.2.1 Exploring the Relationship between Perturbation and Backdoor
where $\delta=\max(\min(\frac{\nabla_{\delta}\ell(f_{\theta}(x),y)}{\|\nabla_{\delta}\ell(f_{\theta}(x),y)\|_{2}},\epsilon),-\epsilon)$ and $\hat{m}$ is a randomly-generated mask of the perturbation. As shown in Fig.4, on both BadNets and Blend attacks, most poisoned samples exhibit good consistency against adversarially generated perturbations $\delta$.
@@
Figure 4: Statistical histogram of training samples (with 5% poison ratio on CIFAR10 using ResNet18).
To further explain this intriguing phenomenon, we propose the corresponding proof based on the recent studies [50] to analyze the characteristics of poisoned samples. We start with the regression solution for the Neural Tangent Kernel (NTK) model as presented in [53, 50, 54]:
$\phi_{t}(\cdot)=\frac{\sum_{i=1}^{N_{b}}Q(\cdot,\boldsymbol{x}_{i})\cdot y_{i}+\sum_{i=1}^{N_{p}}Q(\cdot,\boldsymbol{x}_{i}^{\prime})\cdot y_{t}}{\sum_{i=1}^{N_{b}}Q(\cdot,\boldsymbol{x}_{i})+\sum_{i=1}^{N_{p}}Q(\cdot,\boldsymbol{x}_{i}^{\prime})}$ (4)
Here:
  • $\phi_{t}(\cdot)$ stands for the predictive probability output of the function $f(\cdot;\theta)$ for class $t$,
  • $\boldsymbol{x}_{i}$ are clean training samples,
  • $y_{i}$ are the corresponding one-hot labels (so $y_{i}=1$ if $\boldsymbol{x}_{i}$ belongs to class $t$, and $y_{i}=0$ otherwise),
  • $\boldsymbol{x}_{i}^{\prime}$ are the poisoned samples,
  • $Q(\boldsymbol{x},\boldsymbol{z})=e^{-2\gamma\|\boldsymbol{x}-\boldsymbol{z}\|^{2}}$ is the kernel function with $\gamma>0$,
  • $N_{b}$ and $N_{p}$ denote the number of clean and poisoned samples, respectively.
Assume the target label $y_{t}=1$ and all other labels are 0. Because the training samples are uniformly distributed, $\frac{N_{b}}{k}$ clean samples are labeled as $y_{t}$. This allows us to simplify the equation to:
$\phi_{t}(\cdot)=\frac{\sum_{i=1}^{N_{b}/k}Q(\cdot,\boldsymbol{x}_{i})+\sum_{i=1}^{N_{p}}Q(\cdot,\boldsymbol{x}_{i}^{\prime})}{\sum_{i=1}^{N_{b}}Q(\cdot,\boldsymbol{x}_{i})+\sum_{i=1}^{N_{p}}Q(\cdot,\boldsymbol{x}_{i}^{\prime})}$ (5)
Given that the backdoor sample $\boldsymbol{x}^{\prime}$ generally does not belong to the target class $y_{t}$, the contribution of clean samples in the numerator is negligible compared to the poisoned samples. Therefore, we can approximate $\phi_{t}(\cdot)$ by focusing primarily on the poisoned samples:
$\phi_{t}(\cdot)\geq\frac{\sum_{i=1}^{N_{p}}Q(\cdot,\boldsymbol{x}_{i}^{\prime})}{\sum_{i=1}^{N_{b}}Q(\cdot,\boldsymbol{x}_{i})+\sum_{i=1}^{N_{p}}Q(\cdot,\boldsymbol{x}_{i}^{\prime})}$ (6)
Consider a specific backdoor sample $\boldsymbol{x}^{\prime}$, which is crafted as a mixture of a clean sample $\boldsymbol{x}$ and a target pattern $\boldsymbol{t}$, i.e., $\boldsymbol{x}^{\prime}=(1-m)\odot\boldsymbol{x}+m\odot\boldsymbol{t}$, where $\odot$ denotes element-wise multiplication, and $m$ controls the blending ratio between $\boldsymbol{x}$ and $\boldsymbol{t}$.
When $N_{p}$ approaches $N_{b}$ (i.e., the poisoning ratio is close to 50%), the attack’s efficacy is maximized. In this case, the expression is:
$\phi_{t}(\boldsymbol{x}^{\prime}+\boldsymbol{\delta}^{\prime})\geq\frac{\sum_{i=1}^{N_{p}}Q\left(\boldsymbol{x}^{\prime}+\boldsymbol{\delta}^{\prime},\boldsymbol{x}_{\boldsymbol{i}}^{\prime}\right)}{\sum_{i=1}^{N_{b}}Q\left(\boldsymbol{x}^{\prime}+\boldsymbol{\delta}^{\prime},\boldsymbol{x}_{\boldsymbol{i}}\right)+\sum_{i=1}^{N_{p}}Q\left(\boldsymbol{x}^{\prime}+\boldsymbol{\delta}^{\prime},\boldsymbol{x}_{\boldsymbol{i}}^{\prime}\right)}$ (7)
To get the inequality blow:
$\phi_{t}(\boldsymbol{x}^{\prime}+\boldsymbol{\delta}^{\prime})\geq\frac{1}{2}$ (8)
We need to have:
$\sum_{i=1}^{N_{b}}Q\left(\boldsymbol{x}^{\prime}+\boldsymbol{\delta}^{\prime},\boldsymbol{x}_{\boldsymbol{i}}\right)\leq\sum_{i=1}^{N_{p}}Q\left(\boldsymbol{x}^{\prime}+\boldsymbol{\delta}^{\prime},\boldsymbol{x}_{\boldsymbol{i}}^{\prime}\right)$ (9)
Finally, using the kernel function $Q(\boldsymbol{x},\boldsymbol{z})=e^{-2\gamma\|\boldsymbol{x}-\boldsymbol{z}\|^{2}}$, we see that:
$\sum_{i=1}^{N_{p}}\left(1-e^{-2\gamma\|m\odot(\boldsymbol{t}-\boldsymbol{x}_{i})\|^{2}}\right)>0$ (10)
This inequality shows that the internal term is positive, implying that $f(\boldsymbol{x}^{\prime}+\boldsymbol{\delta}^{\prime})=y_{t}$. Thus, the backdoor attack successfully causes the model to misclassify the poisoned sample as belonging to the target class $y_{t}$, confirming the attack’s effectiveness. Based on the above proof, we can further obtain the theorem below:
Figure 5: Loss landscape of clean sample and poisoned sample.
Theorem 1. Let $\mathcal{D}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{N_{b}}$ be a clean dataset consisting of $N_{b}$ samples, and $\mathcal{D}_{p}=\{(\boldsymbol{x}_{i}^{\prime},y_{t}^{\prime})\}_{i=1}^{N_{p}}$ be a poisoned dataset containing $N_{p}$ samples, both independently and identically distributed (i.i.d.) according to a uniform distribution across $K$ classes. Suppose a deep neural network $f(\cdot;\theta)$ is formulated as a multivariate kernel regression using a Radial Basis Function (RBF) kernel, and shares the same objective as the attackers. For a given attacked sample $\boldsymbol{x}^{\prime}=(\mathbf{1}-\boldsymbol{m})\times\boldsymbol{x}+\boldsymbol{m}\times\boldsymbol{t}$, where $\mathbf{1}$ is a vector of ones, $\boldsymbol{m}$ is a binary mask, $\boldsymbol{x}$ is the original sample, and $\boldsymbol{t}$ is the trigger pattern, we have: $\lim_{N_{p}\to N_{b}}C(\boldsymbol{x}^{\prime}+\delta)=y_{t}$, where $C(\cdot)$ denotes the classification function of the network, $\delta$ represents a perturbation added to the attacked sample, and $y_{t}$ is the target class label.
@@
The theorem above demonstrates that when the size of $\mathcal{D}_{p}$ approaches that of the clean dataset, poisoned samples exhibit robustness against potential perturbations. To gain an intuitive understanding, we visualize the loss landscapes of both clean and poisoned samples (using the Blend attack [25]), as illustrated in Fig. 5. Notably, the loss of a clean sample shows steeper changes when perturbations are applied in the direction of the gradient. This observation further corroborates the previous theoretical analysis.
@@ 4.2.2 Dataset Partition
Algorithm 1 shows the complete process of our dataset partition operation. This algorithm delineates a meticulous process for partitioning a dataset into trusted and potentially compromised subsets by exploiting the variances in the behavior of a model when subjected to perturbed inputs. At its core, the algorithm iterates over each sample $(x,y)$ in the dataset $\mathcal{D}_{u}$, applying a strategically-sized perturbation patch characterized by a radius $r$ and bounded by a perturbation radius $\epsilon$, to the input sample. This is achieved through generating a random patch mask $\hat{m}$ and calculating an optimal adversarial perturbation $\delta$, which is applied to the input $x$ to yield a perturbed counterpart $\hat{x}$. The key metric for partitioning, the KL divergence, is computed between the model’s predictions on the original and perturbed inputs, with these divergence scores being aggregated into a list $Q$. Subsequently, the dataset is partitioned into two subsets, $\mathcal{D}_{u}$ and $\mathcal{D}_{c}$, based on the divergence scores relative to a partition rate $p$, indexing subsets with assumed clean and those potentially poisoned, respectively. This method essentially leverages the sensitivity of backdoored models to perturbations in identifying and isolating suspicious data samples, thus serving as a robust strategy in the preprocessing phase of data handling. With the above steps, we obtain $\mathcal{D}_{u}$ and $\mathcal{D}_{c}$. Here, $\mathcal{D}_{c}$ denotes the selected small number of clean samples. $\mathcal{D}_{u}$ represents the rest of the training data that contains the poisoned samples.
4.2.3 Backdoor Enhancement and Standard Training
@@
4.2.4 Relabel and Relearn
To obtain a completed clean dataset and further improve the performance of the clean model, MeCa relabels the poisoned samples and merges them with clean samples to obtain a clean and complete dataset:
@@ 4.2.4 Relabel and Relearn
5 Experiments
In this section, we experimentally evaluate the performance of the proposed MeCa. Firstly, we show the experimental settings, including the experimental environment, datasets, networks, backdoor attacks, and backdoor defenses. Then, we analyze and summarize the experimental results. Finally, we also present the ablation studies.
5.1 Experimental Settings
5.1.1 Datatsets
We run our experiments on four image datasets, including Imagnette, Tiny ImageNet, CIFAR10, and CIFAR100. The introduction of the above four datasets is shown as follows.
@@ 5.1.1 Datatsets
  • CIFAR10[57]: CIFAR10 dataset is a database of tiny images, containing $50,000$ training images and $10,000$ testing images with $10$ classes.
@@
5.1.2 Networks
We conduct our experiments on three deep learning models: ResNet18, ResNet34, and MobileNetV2. Unless otherwise stated, the default experimental setup assumes the use of ResNet18.
5.1.3 Experimental Environment
All experiments were conducted on a server equipped with Intel i9-9900K, 3.60GHz processor, 32GB RAM, NVIDIA GeForce RTX 3090, and PyTorch.
5.1.4 Backdoor Attacks
We employ $8$ state-of-the-art backdoor attacks to evaluate the proposed MeCa in our experiments, including $3$ dirty-label attacks: BadNet [12], Blend [25], and PhysicalBA [46], one clean-label attack: SIG [48], one input-specific-trigger attack: WaNet [47], and $3$ adaptive attacks: TaCT [43], AdaptiveBlend [49] and AdaptivePatch [49]. The detailed configurations of these backdoor attacks are shown in Table I.
@@ 5.1.4 Backdoor Attacks
5.1.5 Backdoor Defense
We compare MeCa with $7$ established backdoor defense baselines, including FineTuning [58], FinePruning [58], CutMix [59], CLP [24], DBR [19], SCAnFT [43], and ABL [18].
@@ 5.1.5 Backdoor Defense
5.1.6 Evaluation Metrics
We employ the attack success rate (ASR) and accuracy (ACC) to evaluate the performance of the proposed backdoor defense framework. ASR indicates the ratio of the triggered samples that are misclassified as the target label. ACC represents the accuracy of the model on the clean samples. The lower ASR and the higher ACC indicate the better performance of the backdoor defense mechanism.
@@ 5.1.6 Evaluation Metrics
TABLE II: The comparison results (%) between our MeCa and $7$ state-of-the-art backdoor defenses methods against $5$ state-of-the-art backdoor attacks on ResNet18 with $5\%$ poisoned ratio.
No Defense FineTuning FinePruning CutMix CLP DBR SCAnFT ABL MeCa
@@
Imagenette Average 94.14 84.28 68.09 60.72 59.28 60.94 64.76 69.31 23.21 25.76 60.94 28.42 55.13 11.41 47.46 60.49 70.20 0.00
BadNet 92.00 95.20 83.30 96.81 75.05 93.19 81.21 83.30 34.18 0.55 67.47 1.21 69.89 2.09 71.98 0.00 85.72 0.00
@@
CIFAR10 Average 87.12 88.36 83.94 92.61 79.01 78.70 80.57 78.48 32.57 30.81 40.06 41.01 70.22 13.65 82.20 28.72 85.14 0.53
BadNet 72.10 98.20 58.40 91.40 42.00 91.09 42.91 86.03 54.55 83.30 29.35 0.10 56.78 0.00 63.16 0.10 60.53 0.00
@@
CIFAR100 Average 60.62 88.36 61.92 87.21 40.81 57.18 49.15 85.99 45.32 77.99 20.41 52.68 50.12 1.01 62.50 19.15 58.99 0.22
BadNet 37.80 97.95 46.00 98.54 26.87 97.58 31.55 90.84 5.08 84.90 0.70 0.00 32.21 0.00 46.80 0.00 35.73 0.00
@@
TinyImageNet Average 38.79 98.80 48.22 98.55 25.84 83.14 33.95 96.46 9.07 51.90 4.88 46.10 24.61 0.00 38.04 55.67 37.86 0.00
5.2 Comparative Evaluation of MeCa and Existing Defenses
In this paper, we employ $8$ state-of-the-art backdoor attacks (including $5$ familiar backdoor attacks and $3$ adaptive backdoor attacks) to demonstrate the effectiveness of our MeCa. In this section, we take into account $5$ state-of-the-art backdoor attacks and compare the performance of MeCa with $7$ state-of-the-art backdoor defense techniques. Table II shows the performance of the proposed MeCa method on Imagenette, CIFAR10, CIFAR100, and TinyImageNet. Obviously, our MeCa achieves the best results in reducing ASR against most backdoor attacks while maintaining a satisfactory ACC across all $4$ datasets.
Concretely, we mainly have 5 observations from Table II. The detailed analysis is shown as follows:
@@ 5.2 Comparative Evaluation of MeCa and Existing Defenses
(4) All the existing defense methods have an unstable performance against different backdoor attacks. They perform satisfactorily on some backdoor attacks and may be ineffective against the other backdoor attacks. This greatly reduces the availability of these defense methods in real-world scenarios because it is impractical for the defender to know the backdoor methods in advance. For example, ABL performs well on the BadNet attack but has an unsatisfactory performance on the SIG attack. Note that we have observed the different performance of these defense methods in different papers because their hyperparameters are different. We argue that it’s not practical to set different hyperparameters for different attacks because defenders have no knowledge about the potential attacks. Therefore, we use the same hyperparameter configurations for other defenses as well as MeCa.
(5) Compared with the $7$ state-of-the-art backdoor defense techniques, our MeCa can achieve the lowest ASR (the highest ASR is just $2.64\%$) while maintaining a satisfactory performance on the clean samples. Moreover, the proposed MeCa has a more stable performance on different datasets and various backdoor attacks.
TABLE III: The comparison results (%) between our MeCa and $7$ state-of-the-art backdoor defenses methods against $3$ state-of-the-art adaptive attacks on ResNet18 with $5\%$ poisoned ratio.
No Defense FineTuning FinePruning CutMix CLP DBR SCAnFT ABL MeCa
@@
Imagenette Average 94.60 62.83 70.91 4.99 59.00 27.70 68.43 22.53 44.04 10.34 56.88 30.65 54.94 4.43 40.16 32.92 69.99 0.28
TaCT 82.10 63.20 82.63 65.71 75.05 65.71 78.68 41.43 37.91 3.08 77.58 63.19 72.42 62.20 85.16 48.02 82.42 0.00
@@
CIFAR10 Average 83.63 51.27 82.93 60.66 79.45 40.07 71.76 32.34 52.56 3.55 67.69 33.52 74.65 20.81 85.71 50.26 82.86 0.00
TaCT 55.40 22.80 58.70 25.00 42.00 17.71 53.24 17.41 18.72 32.29 41.40 29.15 44.13 0.61 63.26 0.91 61.84 0.00
@@
CIFAR100 Average 47.93 55.47 58.23 56.24 36.47 27.56 46.22 20.82 28.81 13.19 17.14 42.85 52.87 17.81 61.14 12.58 61.61 0.00
TaCT 36.95 13.65 46.55 10.87 26.07 8.05 32.36 19.38 2.62 0.00 0.55 0.00 25.42 0.00 40.71 20.73 38.80 0.00
@@
TinyImageNet Average 36.75 37.13 46.55 61.77 24.43 25.46 32.61 43.10 9.36 9.68 0.40 25.88 27.51 0.00 37.90 57.49 39.78 0.00
5.3 The Resistance to Adaptive Attacks